### v3

# A pedagogical approach to the black hole information issue

Abstract
We provide a pedagogical introduction to the concepts underlying black hole information loss, intended for readers familiar with special relativity and quantum mechanics. We emphasize that there is no paradox of information loss, and that proposals suggesting deviations from well-established theories at arbitrary regimes are inherently contradictory.
Keywords: General relativity. Quantum field theory in curved spacetime. Semiclassical gravity. Hawking radiation. Black hole information issue.
I Introduction
The concept of a region from which nothing can escape due to gravity, known as a black hole, has been a notable front for developments in theoretical physics over the last decades. One of the most distinguished results from these developments is the state of affairs commonly referred to as the black hole information paradox. This unfortunate nomenclature follows mainly from an erroneous interpretation of the implications black holes have for information preservation, which has led to many physically unreliable proposals. In order to fully grasp the nature of this issue and understand associated the misconceptions, it is useful to begin with a review of the current descriptions of the fundamental interactions.
General relativity (GR) Wald1984 is a theory of the gravitational interaction, whose characterization is made by the curvature of spacetime due to the presence of an energy distribution. A qualitative notion of this description is given in fig. 1, inspired by the quote (Misner1973, , p. 5): “Spacetime tells matter how to move; matter tells spacetime how to curve”.
Figure 1: Qualitative description of the gravitational interaction proposed by GR.
This classical [^1] By classical, it is meant that the characterization of the gravitational interaction, given by the spacetime geometry, assumes definite values in any length scale, not only those much larger than atomic ones (i.e., distances of order $10^{-9}\;\text{m}$). theory currently paints the most accurate picture of gravitational phenomena, and thus, is the one we use to describe black holes. On the other hand, the present depiction of the electroweak and strong interactions is given by quantum field theory (QFT) Parker2009 , which can be understood as an extension of quantum mechanics (QM) and special relativity (SR) that treats quantized fields as protagonists of these interactions. Being highly successful in describing observational data and experiments in both of their pertinent regimes of application, these theories can be combined in the mathematically rigorous, but physically incomplete, formalism of quantum field theory in curved spacetime (QFTCS).
The argument behind the physical incompleteness of QFTCS follows from GR, which quantifies how a classical energy distribution, described by the energy-momentum tensor, affects the spacetime geometry. In QFTCS, the quantum field dynamics is governed by the curved spacetime, but one does not know how to generally describe the effects of the quantum field energy content on the spacetime geometry, known as backreaction effects (see fig. 2).
Figure 2: Qualitative description of the interactions in QFTCS. The dashed line represents the fact that in this framework there is no general description of backreaction effects.
Nonetheless, one can rely on physical arguments to qualitatively describe possible backreaction for specific phenomena, which leads to the more complete framework of semiclassical gravity (SG) [^2] Although we distinguish between QFTCS and SG due to the consideration of backreaction effects, some references use these terms as synonyms.. Consequently, following the experimental success of GR and QFT, SG provides the best available description of the fundamental interactions. In other words, SG provides the best formalism available to study quantum effects in and on gravitation.
A notable quantum effect is the Hawking effect, which states that stationary (i.e., time independent) black holes should effectively emit [^3] Hawking radiation does not “get out” from a black hole. an approximate thermal spectrum of particles, known as Hawking radiation. Now, if one considers the implication of a negative energy flux entering the black hole due to local energy-momentum conservation, then the main backreaction effect due to Hawking radiation will be to reduce the black hole mass and other physical parameters, resulting in an effective evaporation process. Hence, after a finite time, a black hole will reach the Plank scale, in the sense that its radius will be of order of the Planck length,
$\ell_{p}=\left(\frac{G\hbar}{c^{3}}\right)^{1/2}\approx 10^{-35}\;\text{m},$ (1)
where $G$ is Newton’s constant, $\hbar$ is Planck’s constant, and $c$ is the speed of light. However, the spacetime curvature at such a regime will be extreme, and dimensional arguments suggest that the smooth spacetime structure postulated by GR should cease to be an adequate description. In this sense, one expects that a full quantum treatment of gravitational degrees of freedom and spacetime structure should be necessary for an accurate picture (i.e., a theory of quantum gravity) at the Planck scale, as semiclassical predictions will no longer be sensible.
To understand how black holes lead to information loss in SG, consider the following line of reasoning. Suppose one starts with a stationary energy distribution, which is described by its mass, angular momentum, electric charge, baryon and lepton number, and any other physical property. From the perspective of QM, complete knowledge of the quantum state of this distribution would correspond to a pure state. An example of this would be a highly diluted cold matter distribution (e.g., a dust cloud) such that one has all the possible information about its configuration. Of course, such a distribution would not be exactly stationary due gravitation and other interactions, but, in distant regions it would be a good approximation to describe it by a stationary spacetime. Now, as this distribution evolves over time, it could form, say, a star. If such a star is massive enough, at the later stages of its life gravitational effects will not be sustained, and a black hole will form as a consequence of gravitational collapse. Even though the black hole formation process is expected to be highly dependent on the energy distribution details and how it collapses, asymmetries of the distribution are expected to be radiated away in the form of gravitational waves. Following a conjecture that black holes must eventually reach a stationary configuration, the initial energy distribution will then give rise to a black hole described only by three parameters: its mass, angular momentum, and electric charge. However, the Hawking effect predicts that such black holes should behave as a gray body, effectively emitting an approximately thermal spectrum of radiation. This radiation carries away the mass, angular momentum, and electric charge of the black hole, and, after a finite time, the black hole might completely evaporate, leaving only radiation. In light of SG, the resulting quantum state would necessarily correspond to a mixed one, which arises in a context in which one does not have all the possible information about the state. Therefore, the complete evaporation of a black hole would take a pure state to a mixed one, corresponding to the fact that most information about the details of the energy distribution that gave rise to the black hole is lost.
Evidently, in order for this conclusion to hold, one would have to trust SG predictions at arbitrary scales. Indeed, even though SG predicts an evaporation process for regimes far from the Planck scale, when a black hole reaches a radius of order $\ell_{p}$, only a theory of quantum gravity would be appropriate to accurately describe its fate. Consequently, it is an interesting open question how a theory of quantum gravity would address black hole evaporation. Nonetheless, the conclusion of information loss in the framework of SG contradicts no known phenomena and established theories; in fact, it is a physical prediction. In contrast, it has been extensively and incorrectly argued in the literature that the conclusion of information loss is paradoxical in light of unitary evolutions in QM. As a result, many inconsistent proposals have come to light to solve a nonexistent paradox.
The purpose of this paper is twofold. First, we provide a pedagogical presentation of the concepts that lead to the conclusion of information loss due to black holes in SG. Second, we show that there is no paradox of information loss, and discuss the flaws in proposals that seek to invalidate SG predictions at arbitrary regimes. The paper is organized as follows. In Sec. II, we discuss black holes in the framework of GR, introducing the necessary ideas for their definition and analysis. In Sec. III, we present the Hawking effect and the implication of quantum entanglement for black holes in SG. In Sec. IV, we show how SG predicts that black holes will result in information loss, and discuss the misconceptions surrounding this conclusion. We conclude in Sec. V with the final remarks. We assume familiarity with SR and QM. The arguments presented are discussed in much more detail in Bergamaschi2024 , which includes an extensive pedagogical review of the technical analysis necessary to support the results stated in this paper. We work with SI units, but we set $c=1$ for all spacetime diagrams simply to allow for more intuitive illustrations. Unless stated otherwise, we will consider four-dimensional spacetimes.
II Classical aspects of black holes
In the purely classical framework of GR, a black hole is defined as a region of spacetime from which nothing can escape. In this section, we will detail the classical concepts necessary to make this notion precise and explain the importance of stationary black holes. Our starting point is to discuss pertinent properties of Minkowski spacetime.
II.1 Minkowski spacetime
In SR, the background for all phenomena is the Minkowski spacetime, which is a collection of events endowed with a four-dimensional measure (i.e., the invariant interval) to every pair of such events given by the Minkowski metric. The concept of causality follows from the possible events that can influence one another, either by a massive particle or an electromagnetic wave. The analysis of which events can be causally connected to one another then defines the causal structure of Minkowski spacetime, which is given by the set of events that can be interpreted as the future and past of events.
For instance, consider an event in Minkowski spacetime, $a$. Associated to $a$ are two sets of events: those that can be connected to it by a timelike or null curve [^4] These curves must obey pertinent differentiability conditions and be oriented in a single direction of time (see (Bergamaschi2024, , sec. 2.2) for more details)., and those that can be connected to it only by a spacelike curve. The latter is the set of events that are causally disconnected from $a$, while the former is the set of events that are causally connected to $a$. In this manner, the notion of future and past for causally connected events arises from the fact that either a massive particle or an electromagnetic wave can be emitted to influence $a$ or to have emanated from $a$. These ideas are illustrated in fig. 3. Notice that any event inside the gray region is causally connected to $a$, as well as those along the null curves that pass through $a$. Events not in the gray regions or on its boundary are causally disconnected from $a$, meaning that only a spacelike curve can connect them to $a$ (i.e., faster than light travel).
Figure 3: Future and past of an event, $a$, in two-dimensional Minkowski spacetime. The illustration contains a timelike curve (blue) and null curves (black).
In particular, the distinction between the future and past sets is given by a predetermined continuous definition of what constitutes going “forwards” or “backwards” in time. In this manner, all curves considered in the following should be understood as “going” from the past to the future.
A helpful way to visualize the entire causal structure of Minkowski spacetime is by a conformal diagram (also known as a Carter-Penrose diagram). Such a diagram allows one to depict the entire spacetime in a finite illustration, while keeping null curves aligned at $\pm 45\degree$. The conformal diagram of Minkowski spacetime is illustrated in fig. 4. In the coordinate system $\{(t,r,\theta,\phi)\}$, where $t$ is the time parameter of static inertial observers in the usual spherical coordinates $\{(r,\theta,\phi)\}$ , the interpretation of the conformal diagram is as follows [^5] In the following, all conformal diagrams will be presented with two angular coordinates suppressed.. The vertical black line is the center of the coordinate system, $r=0$, for different values of $t$. All null curves begin at the three-dimensional null hyperplane $\mathscr{I}^{-}$ and end at three-dimensional null hyperplane $\mathscr{I}^{+}$. Similarly, all timelike curves begin at the point $\mathcal{I}^{-}$ and end at the point $\mathcal{I}^{+}$. Additionally, the red line is a three-dimensional spacelike section (e.g., a $t=\text{constant}$ hyperplane) that begins at $r=0$ and ends at $\mathcal{I}^{0}$. A diagram analogous to that of fig. 3 is also represented in fig. 4, where the null curves emanating from a two-sphere, $\mathscr{A}$, should be understood to extend until they reach $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$. In this manner, the future and past of any other two-sphere can be readily verified by tracing null curves through it, i.e., delimiting its light cone.
Figure 4: Conformal diagram of Minkowski spacetime with two angular coordinates suppressed. The illustration contains null curves (black), a timelike curve (blue), and a three-dimensional spacelike section (red).
One of the most notable features of the conformal diagram of Minkowski spacetime are the two null hyperplanes $\mathscr{I}^{+}$ and $\mathscr{I}^{-}$. These are known as future null infinity and past null infinity, respectively. They can be interpreted as “infinitely distant regions” of Minkowski spacetime such that only light rays can reach, to the future and to the past. This is due to the fact that all null curves begin at $\mathscr{I}^{-}$ and end at $\mathscr{I}^{+}$. To visualize this, consider a light ray that was emitted far away from the origin of the coordinate system (i.e., $r\gg 0$) at early times (i.e., $t\ll 0$). If the light ray was emitted so that its radial coordinate decreases over time (i.e., dr/dt=cdr/dt=-c), then it will eventually reach the origin of the coordinate system, $r=0$. After that, its radial coordinate will increase over time (i.e., $dr/dt=c$), and it will eventually reach the regions far away from the origin at late times (i.e., $t\gg 0$). The regions $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$ appear in this line of reasoning when one takes the limit of $r\to\infty$ (i.e., far away region) and $t\to\pm\infty$ (i.e., early and late times). In particular, the process of the light ray reaching $r=0$ and thus being described by a different null curve is illustrated by the “reflection” of the null curve in fig. 4 (see also fig. 5 for more examples).
To conclude our analysis of Minkowski spacetime, we mention an important property of its causal structure. From its conformal diagram, it can be deduced that any three-dimensional spacelike section that begins at $r=0$ and extends to $\mathcal{I}^{0}$ will be intersected exactly once by all null curves. More specifically, consider two three-dimensional spacelike sections that begin at $r=0$ and extend to $\mathcal{I}^{0}$, denoted by $\Sigma_{0}$ and $\Sigma_{1}$. From fig. 5, it is possible to see that all null curves emerge from $\mathscr{I}^{-}$ and intersect $\Sigma_{0}$ and $\Sigma_{1}$ exactly once, either on their way to reach $r=0$ or on their way to reach $\mathscr{I}^{+}$ after “reflecting” at $r=0$. Because of this property, $\Sigma_{0}$ and $\Sigma_{1}$ are referred to as Cauchy surfaces. Spacetimes possessing a Cauchy surface are said to be globally hyperbolic, which can be interpreted as those that obey pertinent conditions on causality and determinism, allowing one to formulate a well-posed problem for dynamics.
Figure 5: Conformal diagram of Minkowski spacetime illustrating two Cauchy surfaces (red), $\Sigma_{0}$ and $\Sigma_{1}$.
Additionally, globally hyperbolic spacetimes can be described in terms of a foliation of Cauchy surfaces, i.e., they can be written as a collection of Cauchy surfaces. An example of this is a foliation of Minkowski spacetime in terms of $t=\text{constant}$ hyperplanes. Although a foliation is not unique, in a given foliation each Cauchy surface can be interpreted as an “instant of time” of the spacetime. In this sense, information on a Cauchy surface can be used to determine the entire development of the spacetime, i.e., physical conditions on any other Cauchy surface. Indeed, because each null curve must intercept a Cauchy surface exactly once, information can be interpreted to be carried over (i.e., conserved) from one Cauchy surface to the others. This can be exemplified by considering both $\Sigma_{0}$ and $\Sigma_{1}$, as all null curves that intersect $\Sigma_{0}$ will eventually intersect $\Sigma_{1}$. Conversely, complete predictability from physical conditions at some “instant of time” will not occur for spacetimes that fail to be globally hyperbolic. Consequently, it is challenging to envision how information could be conserved when one is analyzing non-globally hyperbolic spacetimes.
II.2 Definition
In GR, the flat geometry of Minkowski spacetime is merely a particular solution for the spacetime geometry when no energy distribution is present. Following the description of fig. 1, spacetime is now postulated to curve due to the presence of an energy distribution. Nevertheless, the notion of a causal structure is still pertinent to a curved spacetime, in the sense that one can rely on the same arguments as those presented in Minkowski spacetime to define the future and past of events. Our focus now is to use these ideas to give a rigorous definition of a black hole.
The intuitive notion of a black hole is a region from which nothing can escape. To make this notion precise, one needs to describe what it means “to escape from” and to where it could be pertinent to “escape”. Notably, the idea that there is a restriction on the behavior of timelike and null curves is precisely captured by the concept of causally disconnected events. Namely, if there is no null or timelike curve connecting two events, then they are causally disconnected. Consequently, “to escape from” can be understood as the possibility of events to be causally connected. Additionally, the idea of where null curves could contemplate “escaping to” is given by future null infinity, $\mathscr{I}^{+}$. As discussed in Sec. II.1, in Minkowski spacetime all null curves end at $\mathscr{I}^{+}$, which means that, naturally, one expects that null curves would eventually reach $\mathscr{I}^{+}$, “escaping from” the rest of spacetime. In this sense, the definition of a black hole can be made precise in spacetimes possessing a region similar to $\mathscr{I}^{+}$. These are known as asymptotically flat spacetimes, as the “infinitely distant regions” have a Minkowski-like character. In other words, the “infinitely distant regions” to the past and the future in asymptotically flat spacetimes behave similarly to $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$.
In this manner, the black hole region of an asymptotically flat spacetime is defined as the set of events that are causally disconnected from $\mathscr{I}^{+}$. Physically, this can be interpreted as the fact that no electromagnetic wave emanating from regions inside the black hole can reach $\mathscr{I}^{+}$. In other words, nothing can escape from the black hole due to gravity, not even light. An example of this is given in fig. 6, which illustrates the conformal diagram of a spherically symmetric and electric uncharged black hole, known as a Schwarzschild black hole. The spacetime describing such a black hole is known as Schwarzschild spacetime, and the interpretation of its diagram is the following. Region II, highlighted in blue, is the black hole region. Region I is the exterior vacuum region, i.e., no classical energy distribution is present. Region III is the interior region of the spherically symmetric energy distribution that results in the black hole. Region II is delimited by the radial coordinate known as Schwarzschild radius,
$r_{s}=\frac{2GM}{c^{2}}\approx 10^{3}\left(\frac{M}{M_{\odot}}\right)\;\text{m},$ (2)
where $M_{\odot}$ is the Sun’s mass. The boundary of the black hole region, in this case characterized by the null curve obeying $r=r_{s}$, is known as the event horizon. Additionally, the thick white line at the top of the diagram represents the singularity inside the Schwarzschild black hole, a “pathological” region of the spacetime. The rest of the diagram boundary has the same interpretation as that of the conformal diagram of Minkowski spacetime.
Figure 6: Conformal diagram of Schwarzschild spacetime illustrating the exterior vacuum region (I), the black hole region (II) (highlighted in blue), and the collapsing energy distribution (III).
To visualize why region II is the black hole region, consider some examples of null and timelike curves, as shown in fig. 7.
Figure 7: Conformal diagram of Schwarzschild spacetime illustrating pairs of timelike curves (blue) and null curves (black).
The timelike curve $\lambda_{2}$ is an example of the world-line of a massive particle that stays outside region II and that will end at $\mathcal{I}^{+}$. In contrast, the timelike curve $\lambda_{1}$ is an example of the world-line of a massive particle that enters region II, and its world-line ends not in $\mathcal{I}^{+}$, but rather, in the singularity. Similarly, the null curve $\gamma_{2}$ reaches $\mathscr{I}^{+}$, while $\gamma_{1}$ enters region II and eventually reaches the singularity. Indeed, any event in region II will not only not be able to communicate with events outside of it, but any massive particle or electromagnetic wave emanating from it will unavoidably reach the singularity. Notice that this is also the case for the energy distribution in region III. More precisely, after it collapses (i.e., reaches $r_{s}$), all the distribution will eventually reach the singularity.
In this manner, defining the black hole region as the set of events causally disconnected from $\mathscr{I}^{+}$ captures the idea of a region from which nothing can escape. However, the fact that $\mathscr{I}^{+}$ represents “infinitely distant regions” means that one would require complete knowledge of spacetime in order to analyze which events are causally connected to it. Due to this, it is convenient to restrict the definition of black holes to spacetimes that are globally hyperbolic at least outside the black hole and on the event horizon. With this, one would be able to study dynamical aspects of black holes in a covariant manner using Cauchy surfaces. Indeed, from figs. 8 and 9, it is easy to see that Schwarzschild spacetime is globally hyperbolic by using the same criterion as for Minkowski spacetime (see fig. 5). On the other hand, without this condition of global hyperbolicity, it is difficult to perceive a way to discuss a physically significant notion of time evolution of black holes.
Figure 8: Conformal diagram of Schwarzschild spacetime illustrating three Cauchy surfaces (red), $\Sigma_{0}$, $\Sigma_{1}$, and $\Sigma_{2}$.
Figure 9: Sketch of the Finkelstein diagram corresponding to the conformal diagram shown in fig. 8. In this diagram, only null curves whose radial coordinate decrease over time are at $\pm 45\degree$, as shown by the “tilting” of the light cones as they approach $r=r_{s}$.
II.3 Stationary black holes
Stationary black holes are those that do not change over time. Their analysis is pertinent mainly because of two reasons. First, one expects that all physical black holes (i.e., those resulting from gravitational collapse) [^6] In the following, we will only consider physical black holes. should be eventually described, to a good order of approximation, by a time independent configuration. Second, theorems in GR state that any stationary black hole can be described by only three parameters.
The first reason comes from the expectations regarding the product of gravitational collapse that results in a black hole. In particular, although details of the gravitational collapse can greatly affect the spacetime geometry, at sufficiently “late times” after its formation (as “measured” by Cauchy surfaces), a black hole is expected to reach a stationary configuration due to interaction with other sources of energy and the emission of gravitational waves (i.e., undulations of curvature propagating in spacetime). This conjecture is corroborated by the behavior of other systems, e.g., electromagnetic ones, in which a time dependent distribution of electric charges radiates away the higher order multipole moments and eventually “settles down” to a stationary distribution.
The second reason comes from the results known as the black hole uniqueness theorems. These theorems show that any stationary black hole must be described uniquely by its mass, $M$, angular momentum, $L$, and electric charge, $q$. The most general stationary black hole solution is known as a Kerr-Newman black hole described by the parameters $(M,L,q)$. Evidently, a Schwarzschild black hole is a nonrotating (i.e., $L=0$) and electric uncharged (i.e., $q=0$) Kerr-Newman black hole. Coupled with the conjecture that black holes should eventually reach a stationary configuration, the black hole uniqueness theorems leads to the conclusion that any black hole will eventually have only three degrees of freedom.
III Semiclassical aspects of black holes
In this section, we will discuss the two main quantum effects that are pertinent to the conclusion of information loss in black holes. The first concerns the effective particle creation by black holes, known as the Hawking effect. The second is about an implication of entanglement between regions of spacetime in the framework of SG. Our starting point is to explain the concept of a particle content in QFT.
III.1 Hawking effect
In QFT, a quantized field is the protagonist, being described by a collection of quantum harmonic oscillators in the spacetime. In this framework, what one would classically refer to as particles are now viewed as quanta or excitations of the oscillators describing the quantum field. As one would proceed in QM, the evaluation of a quantum-state particle content is done by the computation of the expectation value of the number operator. Consequently, dynamical processes that generate excitations of the quantum field can be understood as resulting in particle creation. This is the fundamental concept behind the Hawking effect.
In essence, an observer “near” $\mathscr{I}^{-}$ would identify an initial no particle state (i.e., one with no quanta or excitations) as the vacuum state, $|0\rangle_{\mathscr{I}^{-}}$. An example of such an observer is one at the early stages of its life, with a world-line given by $\lambda_{1}$ or $\lambda_{2}$ in fig. 7. Now, let $|0\rangle_{\mathscr{I}^{+}}$ denote the vacuum state for an observer “near” $\mathscr{I}^{+}$. An example of such an observer is one that remains outside the black hole at the late stages of its life, with a world-line also given by $\lambda_{2}$ in fig. 7. Notably, due to gravitational collapse and black hole formation, one finds that $|0\rangle_{\mathscr{I}^{-}}$ does not simply evolve to $|0\rangle_{\mathscr{I}^{+}}$, as illustrated in fig. 10.
Figure 10: Creation of particles due to the process of gravitational collapse and black hole formation Fabbri2005 .
This means that observers at $\mathscr{I}^{+}$ perceive particles.
The content of the particle spectrum measured at $\mathscr{I}^{+}$ can be evaluated to be predominantly proportional to that of a black body, whose temperature is given by the Hawking temperature,
$T=\frac{\hbar\kappa}{2\pi ck_{B}},$ (3)
where $k_{B}$ is Boltzmann’s constant and $\kappa$ is the black hole surface gravity. This result can be interpreted as stating that a Kerr-Newman black hole will effectively behave predominantly as a gray body [^7] The more detailed form of the spectrum contains gray body factors due to transmission probabilities., whose temperature is proportional to its surface gravity [^8] Although the local acceleration necessary to keep a massive particle still at the event horizon is divergent, the surface gravity is a finite quantity. This is because it is defined as the acceleration as measured from $\mathscr{I}^{+}$.. This effective particle creation by black holes is known as the Hawking effect, while the approximately thermal radiation predicted by it is referred to as Hawking radiation. As expected from the uniqueness theorems, a detailed evaluation of the spectrum for a Kerr-Newman black hole reveals that the emission of particles depends only on the parameters $(M,L,q)$. Also, the emission of particles with angular momentum and electric charge with the same sign as those of the black hole have a higher probability than those with opposite sign. Finally, the Hawking temperature of a Schwarzschild black hole is
$T\approx 6.18\;10^{-8}\left(\frac{M}{M_{\odot}}\right)^{-1}\;\text{K}.$ (4)
III.2 Entanglement
Entanglement is a feature that arises in composite quantum systems, limiting one’s ability to obtain complete information about an individual system. Namely, suppose one is dealing with the two-particle state
$|\psi\rangle=\frac{1}{\sqrt{2}}|\uparrow_{1}\rangle\otimes|\downarrow_{2}\rangle+\frac{1}{\sqrt{2}}|\downarrow_{1}\rangle\otimes|\uparrow_{2}\rangle.$ (5)
Evidently, this is a pure composite state, because it is a product state of system $1$ (corresponding to particle 1) and system $2$ (corresponding to particle 2) that is represented by a single ket in the product Hilbert space. The fact that it is a pure state means that one has all the possible information about the composite system. However, by ignoring the degrees of freedom of system $1$ (i.e., by “tracing-out” its degrees of freedom in the state $|\psi\rangle$) one finds that system $2$ is described by the density operator
$\hat{\rho}=\frac{1}{2}|\uparrow_{2}\rangle\langle\uparrow_{2}|+\frac{1}{2}|\downarrow_{2}\rangle\langle\downarrow_{2}|,$ (6)
i.e., a mixed state. This means that, by restricting one’s attention only to system $2$, there is now a probability distribution of states, and one cannot know for certain what state describes the individual system. It is in this sense that entanglement affects how much information one can acquire about the state of a single system which is a part of a larger one.
In the context of SG, entanglement is an essential feature in light of what one would expect to be a physically acceptable class of states. In particular, since one is interested in describing backreaction effects (see fig. 2), it is of significance to understand properties that a state should posses in order to make sense of a notion of energy content. To understand this in more detail, consider the following line of reasoning.
Recall that the ground state for a quantum harmonic oscillator has energy $\hbar\omega/2$. If one considers the infinite collection of quantum harmonic oscillators that describe a quantum field, then an associated vacuum state has a divergent energy content. Indeed, the nontrivial structure of the vacuum state in QFT leads to several effects that have been experimentally observed. However, no problems arise in the theoretical framework due to procedures that deal with the divergent quantities that appear (i.e., regularization and renormalization procedures). For example, in Minkowski spacetime, one performs a vacuum energy subtraction procedure using the natural vacuum state defined in it, the Minkowski vacuum, $|0\rangle_{M}$. This procedure allows one to give a finite notion of the energy content of a state, at least for states belonging to a certain class.
Yet, such a procedure is not valid for general curved spacetime because there is no analog of the Minkowski vacuum. Nevertheless, one can select a class of states that satisfy the condition of having a divergent character similar to that of the Minkowski vacuum. States obeying this property are known as Hadamard states, and it can be argued that in order for a state to be physically acceptable it must be a Hadamard state. Indeed, it is possible to define a procedure that results in a finite energy content for Hadamard states Wald1994 ; Parker2009 .
One of the main features of Hadamard states is that they imply entanglement over spacelike separated regions. In fig 11, we illustrate how this is of significance for black holes using two Cauchy surfaces, $\Sigma_{0}$ and $\Sigma_{1}$. Assuming that the initial state is a pure Hadamard state, then the field state at $\Sigma_{0}$ is also pure. The dynamical evolution of physical conditions from $\Sigma_{0}$ to any other Cauchy hypersurface will be unitary, and thus, will maintain its purity. For instance, for the “instant of time” $\Sigma_{1}$, the total field state is also pure. But, because $\Sigma_{1}$ intersects the black hole region, observers outside the black hole will have access only to a portion of $\Sigma_{1}$ (the dark red portion). Consequently, due to the entanglement implied by the Hadamard state, the field state inside the black hole “at a time” $\Sigma_{1}$ (the purple portion) will be entangled with the one outside. Hence, even though the field state at $\Sigma_{1}$ is pure, observers outside the black hole will not be able to acquire complete information about the state in the dark red portion. Obviously, they can always blame the black hole for concealing information about the quantum field, in the sense that information about the state inside the black hole could be accessed once one enters it. Nevertheless, information about the state inside the black hole will never be able to be shared with observers outside the black hole.
Figure 11: Conformal diagram of a Schwarzschild spacetime illustrating two Cauchy hypersurfaces, $\Sigma_{0}$ (red) and $\Sigma_{1}$ (divided in two portions, dark red and purple).
IV Time evolution of semiclassical black holes
In this section we show how the semiclassical properties of black holes discussed so far leads to the conclusion of information loss. We will show why this conclusion is in agreement with semiclassical arguments, and does not earn the adjective “paradoxical”. Additionally, we will present misconceptions surrounding some proposals that seek to change the conclusion of information loss.
IV.1 Information loss
Although calculating general backreaction effects is not a trivial task for four-dimensional spacetimes, one expects that the main consequence of the Hawking radiation is to reduce the parameters of the black hole at a given time, i.e., its mass, angular momentum, and electric charge. Consequently, one would have an effective evaporation process. This qualitative description of a possible backreaction effect is corroborated by physical arguments. More precisely, the energy reduction is merely a consequence of the fact that the positive energy flux to $\mathscr{I}^{+}$ implies that there must exist a negative energy flux going into the black hole, while the reduction of the other parameters is due to the fact that the spectrum of emitted particles tends (i.e., is more likely) to carry away angular momentum and electric charge. Because of the characteristics of the spectrum, a Kerr-Newman black hole will quickly (in comparison with time scales of interest) become a Schwarzschild black hole. Therefore, for simplicity, our analysis will be focused on a Schwarzschild black hole.
As per equation (4), the Hawking temperature of a Schwarzschild black hole is inversely proportional to its mass, which means that the typical energy of created particles increases as a Schwarzschild black hole evaporates. In particular, this relation between particle energy spectrum and black hole mass ensures that created particles will have a small influence on the spacetime metric for most of the evaporation process, so that neglecting additional backreaction effects will be a good approximation when $r_{s}\gg\ell_{p}$ (i.e., far from the Planck scale). In this sense, details of backreaction effects will pose little influence over the event horizon and the exterior region. Consequently, the spacetime geometry can be described by a sequence of quasi-static processes in which the black hole radius decreases slowly, with the process of energy loss being approximated by Stefan’s law. This argument leads one to the conclusion that a black hole should completely evaporate, i.e., radiate away its mass, in a finite amount of time (as “measured” by Cauchy hypersurfaces). Evidently, at some point in time $r_{s}\sim\ell_{p}$, hence, complete evaporation is only a valid prediction if one considers that no deviations from SG occur at that scale. For the moment, let us consider that to be the case and analyze the consequences of a complete evaporation process.
The conformal diagram of a Schwarzschild black hole that resulted from a spherically symmetric collapse and evaporated completely is illustrated in fig. 12. Notably, the spacetime fails to be globally hyperbolic.
Figure 12: Conformal diagram of a Schwarzschild black hole that formed from a spherically symmetric distribution of energy and evaporated completely.
This can be verified by the fact that some null curves do not intersect $\Sigma_{2}$. More precisely, although after complete evaporation events at $r=0$ are again causally connected to $\mathscr{I}^{+}$, there are null curves that never intersect $\Sigma_{2}$ because they reach the singularity [^9] Notice that the event of complete evaporation corresponds to a singularity that is causally connected to $\mathscr{I}^{+}$. Thus, null curves emanating from it do not intersect any section such as $\Sigma_{0}$ or $\Sigma_{1}$.. Thus, physical conditions at $\Sigma_{2}$ will not suffice to determine those at $\Sigma_{0}$ or $\Sigma_{1}$. Fundamentally, information loss can be attributed to the fact that some null curves reach the singularity. Indeed, information loss is also evident from the fact that the field state at a spacelike section such as $\Sigma_{2}$ (“after” complete evaporation) is not sufficient to determine the complete field state at either $\Sigma_{0}$ or $\Sigma_{1}$. In particular, considering the same color scheme of fig. 11, at $\Sigma_{0}$ one has an initial pure Hadamard state. At $\Sigma_{1}$, the total field state is still pure, but the field state outside the black hole is a mixed one due to entanglement with the black hole interior. Consequently, after complete evaporation, the field state at any spacelike section such as $\Sigma_{2}$ will be a mixed one. In essence, there is still entanglement with black hole interior, even though it “no longer” exists. It is in this sense that black holes will result in a dynamical evolution from a pure state to a mixed one (i.e., a non-unitary dynamical evolution).
The generality of this analysis follows from the arguments presented in Secs. II and III. Namely, suppose one starts with an isolated energy distribution described by a pure Hadamard state. If this distribution is massive enough, gravitational effects will not be sustained, and a black hole will form. Following the emission of gravitational waves, the black hole will eventually be described to a good order of approximation by a Kerr-Newman configuration. The Hawking effect together with the qualitative description of backreaction means that the Kerr-Newman black hole will (quickly) result in a Schwarzschild black hole. One is then led to the same line of reasoning presented above, where information about the field state after complete evaporation does not suffice to determine the initial state.
Now, the expectation of information conservation is only justified in a globally hyperbolic spacetime, where information is carried from one Cauchy surface to another. In fact, QFT predicts that information on a Cauchy surface will not in general be completely deposited in a future region that fails to be a Cauchy surface. This means that even in a globally hyperbolic spacetime one can still find examples of “information loss”. For example, in Minkowski spacetime one can consider the field evolution from a Cauchy surface, $\Sigma_{0}$, to an asymptotically null hyperboloid, $\Sigma_{1}$ (see fig. 13) Wald1994 .
Figure 13: Spacetime diagram of Minkowski spacetime illustrating a Cauchy hypersurface, $\Sigma_{0}$, and an asymptotically null hyperboloid, $\Sigma_{1}$ Wald1994 .
Because there are null curves that do not intersect $\Sigma_{1}$, a pure state at $\Sigma_{0}$ would necessarily evolve to a mixed one at $\Sigma_{1}$. As such, dynamical evolution from $\Sigma_{0}$ to $\Sigma_{1}$ would not be unitary as it does not constitute a closed system (i.e., an isolated system). Evidently, this is not at odds with the postulates of QM, which requires unitary evolutions only for closed systems.
Similarly, information loss due to the formation and complete evaporation of black holes can be interpreted as a consequence of the singularity resulting in an open system (i.e., a non-isolated system). In this sense, the prediction of information loss in SG is a physical one, and hence, should not be interpreted as paradoxical.
IV.2 Misconceptions about alternatives to information loss
It has been extensively advocated in the literature that such pure-to-mixed-state evolution implies a violation of physical principles and conjectures regarding black holes (e.g., that black holes are ordinary thermodynamic systems Almheiri2021 ), and therefore, black hole evaporation must be a unitary process. The incorrect line of reasoning that leads to these interpretations views information loss as paradoxical, and refers to the state of affairs surrounding the product of black hole evaporation as the black hole information paradox.
Notably, if one proposes a different description of black hole evaporation that leads to information conservation, it is imperative to preserve semiclassical physics far from the Planck scale. More precisely, a proposal that seeks to change semiclassical predictions at arbitrary regimes is inherently contradictory. Examples of such contradictory proposals are those that support the view that black holes never form (e.g., Mathur2005 ) and those arguing that Hawking radiation should strongly deviate from thermality (e.g., Almheiri2021 ). Notice that formation of black holes and the derivation of the predominant thermal character of the Hawking radiation are robust results that do not rely on specific regimes. For instance, consider the curvature scales at the event horizon of a Schwarzschild black hole, which are proportional to $r_{s}^{-2}$. From eq. 2, in order to avoid the formation of any Schwarzschild black hole, one would have to find deviations from SG predictions in arbitrary curvature scales. The exact same argument applies to deviations from thermality by Hawking radiation. That is, a strong deviation from thermality at arbitrary regimes could only be accounted for by strong deviations from the semiclassical picture far from the Planck scale. However, QFT and GR are well-tested at low curvature regimes (e.g., Solar system-like curvature settings), and, because no deviations are found, relying on arguments that would require such descriptions to fail at arbitrary regimes is contradictory. On the other hand, nothing can be said about proposals that change the product of the evaporation process only at the Planck scale, neither in favor nor against them.
V Conclusions
In this paper, we reviewed the classical and semiclassical description of black holes and presented the conjectures and hypotheses that lead to the conclusion of information loss. We show that the physical non-unitary dynamical evolution that follows from black hole complete evaporation can be interpreted as a consequence of the singularity “opening” the system. Consequently, the conclusion of information loss in SG is not paradoxical. Additionally, we discussed the contradictory nature of proposals that suggest deviations from SG at arbitrary scales.
While acknowledging that a theory of quantum gravity is required to make precise predictions about the outcome of black hole evaporation, particularly regarding backreaction effects and Planck scale physics, we emphasize that the hypotheses, conjectures, and approximations of SG are well justified. Nonetheless, it remains a compelling open issue how black hole evaporation will be addressed within an adequate theory of quantum gravity, and what novel physics it could provide.
Acknowledgments
The author thanks Esmerindo de Sousa Bernardes for his support during the early stages of this work, Daniel A. T. Vanzella for extensive support and for comments on a draft, and George E. A. Matsas for many conversations on black hole physics and support during the later stages of this work. The author also thanks two anonymous referees for detailed and helpful comments. This study was financed by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001.